# Artificial Intelligence for Microbiology and Microbiome Research

### v1

@@ 3 Application Scenarios
There are numerous applications of AI techniques in microbiome research. We can briefly group those applications into the following scenarios: taxonomic profiling, functional annotation & prediction, microbe-X interactions, microbial ecology, metabolic modeling, precision nutrition, clinical microbiology, prevention & therapeutics. For each application scenario, there are many specific tasks. In the following, we will present each of the specific tasks and the representative AI methods.
@@ 3.1 Taxonomic Profiling
Figure 2: Application scenarios of AI in microbiology and microbiome research.
A fundamental goal of microbiology and microbiome research is determining the compositions of microbial communities, i.e., identifying and quantifying different types of microorganisms (such as bacteria, fungi, viruses, and archaea) present in a given sample. This involves analyzing their relative abundances and diversity, often using DNA sequencing techniques. Currently, three generations of DNA sequencing techniques are available for microbiome research. The first-generation sequencing utilizes the chain termination method, offering read lengths of 500-1000 base pairs [76]. Second-generation sequencing, also known as next-generation sequencing (NGS), includes methods such as pyrosequencing, sequencing by synthesis, and sequencing by ligation, with read lengths ranging between 50 and 500 bp [77]. Two key NGS applications in microbiome research are (1) amplicon sequencing, which targets small fragments of one or two hypervariable regions of the 16S rRNA gene (for archaea and bacteria) or 18S rRNA gene (for fungi); and (2) metagenomic shotgun sequencing, which comprehensively samples all genes in all organisms present in a given community. NGS also offers short reads, with read lengths reaching 50-500 bp [78, 79, 77]. The third-generation sequencing performs single-molecule sequencing, offering long reads with lengths reaching tens of kilobases on average [80]. In the following, we discuss applications of AI techniques in various aspects of taxonomic profiling.
@@ 3.1.1 Metagenome assembly
Metagenomics refers to the direct study of the entire genomic information contained in a microbial community. Metagenomics avoids isolating and culturing individual microorganisms in a community and provides a way to study microorganisms that cannot be isolated and cultured. There are two main approaches for processing metagenomic sequencing data: (1) assembly-based and (2) reference database-based. The goal of the assembly-based approach is to construct and annotate the so-called metagenome-assembled genomes (MAGs) [81]. The construction and annotation of MAGs have greatly promoted our understanding of microbial populations and their interactions with the environment. It is worth noting that most MAGs represent new species, which helps to understand the so-called microbial dark matter. The process of constructing MAGs includes two main steps: assembly and binning. Assembly refers to the process of reconstructing longer sequences (contigs) from short DNA reads obtained through sequencing. This involves piecing together overlapping reads to form continuous sequences that represent parts of the genomes present in the microbial community.
@@
Deep learning has been widely used in the quality control of metagenomic assembly. Many factors (e.g., sequencing errors, variable coverage, repetitive genomic regions, etc.) can produce misassemblies. For taxonomically novel genomic data, detecting misassemblies is very challenging due to the lack of closely related reference genomes. Deep learning methods can identify misassembled contigs in a reference-free manner. Representative methods include DeepMAsEd [82] and ResMiCo [83]. DeepMAsEd is based on CNN. Denote a contig as a sequence of nucleotides. At each position in the sequence, the concatenation of two types of information (raw sequence and read-count features) yields the input vector. To train and evaluate DeepMAsEd, one can generate a synthetic dataset of contigs, read counts, and binary assembly quality labels. As an extension of DeepMAsEd, ResMiCo is based on ResNet, a variant of CNN. The key feature of ResNet is the introduction of skip connections, which effectively solves the degradation problem of deep neural networks [84]. Compared to DeepMAsEd, ResMiCo leveraged a much more informative input vector computed from raw reads and contigs. Moreover, ResMiCo was trained on a very large and varied dataset. Through thorough validation, it was demonstrated that ResMiCo significantly outperforms other methods in accuracy, and the model remains robust when faced with novel taxonomic diversity and different assembly methods. We notice that both DeepMAsEd and ResMiCo used a carefully designed input vector. It would be interesting to explore if we can use a more advanced deep learning architecture (e.g., the Transformer) or a hybrid learning approach (e.g., CNN + RNN) to directly deal with the raw sequence data, avoiding the manual design of the input vector.
@@ 3.1.2 Metagenome binning
Metagenomic binning involves grouping those assembled sequences into clusters (bins or MAGs) that correspond to different species or genomes. Metagenomic binning helps in identifying and categorizing the different microorganisms present in a metagenomic sample, even if they are not fully assembled into complete genomes. There are many methods for metagenomic binning [85, 86, 87, 88]. Several binning methods are based on deep learning, e.g., VAMB [89], CLMB [90], SemiBin [91], GraphMB [92], and COMEBin [93]. VAMB (Variational Autoencoders for Metagenomic Binning) uses VAE to encode sequence coabundance and $k$-mer distribution information, and clusters the resulting latent representation into genome clusters and sample-specific bins [89]. As an extension of VAMB, CLMB (Contrastive Learning framework for Metagenome Binning) can efficiently eliminate the disturbance of noise and produce more stable and robust results [90]. CLMB is based on contrastive learning, an machine learning approach that focuses on extracting meaningful representations by contrasting positive and negative instances [90]. SemiBin employs deep siamese neural networks to exploit the information in reference genomes, while retaining the capability of reconstructing high-quality bins that are outside the reference dataset [91]. Here, a siamese neural network (a.k.a. twin neural network) is a neural network that uses the same weights while working in tandem on two different input vectors to compute comparable output vectors [94]. GraphMB integrates GCN with assembly graphs to improve binning accuracy [92]. It models each contig using VAE for feature generation and aggregates these features using a GCN. This method accounts for read coverage in its loss function and uses iterative medoid clustering to finalize the binning. COMEBin is the latest metagenomic binning method [93]. This method is based on contrastive multiview representation learning. It introduces a data augmentation approach that generates multiple views for each contig, enabling contrastive learning and yielding high-quality representations of the heterogeneous features. Moreover, it incorporates a “Coverage module” to obtain fixed-dimensional coverage embeddings, which enhances its performance across datasets with varying numbers of sequencing samples. It also adapts an advanced community detection algorithm, Leiden, specifically for the binning task, considering single-copy gene information and contig length. COMEBin outperformed VAME and SemiBin on various simulated and real datasets, especially in recovering near-complete genomes from real environmental samples.
@@ 3.1.4 Nanopore sequencing basecalling
Nanopore sequencing technology has enabled inexpensive long-read sequencing with reads longer than a few thousand bases [115]. The basic principle of nanopore sequencing is to pass an ionic current through a nanopore and measure the change in current when a biomolecule passes through or approaches the nanopore. Information about the change in current can be used to identify the molecule, a process often referred to as basecalling. There are two challenges in basecalling. First, the current signal level is most dominantly influenced by the several nucleotides that reside inside the pore at any given time, rather than a single base. Second, DNA molecules do not translocate at a constant speed. Basecalling is conceptually similar to speech recognition. Both processes involve interpreting complex signals to extract meaningful sequences—DNA bases in the case of basecalling, and spoken words in the case of speech recognition. Much like the evolution of speech recognition methods, computational methods for basecalling have evolved from statistical tests to hidden Markov models and finally deep learning models. Those methods are often referred to as basecallers.
Various deep learning models have been developed for basecalling. Chiron is the first deep learning model that can translate raw electrical signal directly to nucleotide sequence [116]. It applied a CNN to extract features from the raw signal, an RNN to relate such features in a temporal manner, and a connectionist temporal classification (CTC) decoder to create the nucleotide sequence. Here, CTC enabled us to generate a variant length base sequence for a fixed-length signal window through output-space searching, avoiding explicit segmentation for basecalling from raw signals. Similar to the Chiron architecture, SACall [117] (CATCaller [118] or Bonito [119]) integrated CNN with Transformer (Lite Transformer or LSTM) and CTC. Mincall [120] (or Causalcall [121]) directly integrated ResNet (or causal dilated CNN) with CTC. Halcyon used a different architecture. It combines a novel inception-block-based CNN module, an LSTM-based encoder, and an LSTM-based decoder using an attention mechanism. The inception-block-based CNN module aims to extract local features of input raw signal and reduce the dimension of the input timestep axis. The LSTM-based encoder captures long-time dependencies in the timestep dimension and deals with the variable lengths of inputs. The attention mechanism allows the decoder to focus on specific parts of the input sequence when generating each element of the output sequence.

### v2

@@ 3 Application Scenarios
There are numerous applications of AI techniques in microbiology/microbiome research. We can briefly group those applications into the following scenarios: taxonomic profiling, functional annotation & prediction, microbe-X interactions, microbial ecology, metabolic modeling, precision nutrition, clinical microbiology, prevention & therapeutics (Fig.1). For each application scenario, there are many specific tasks. In the following, we will present each of the specific tasks and the representative AI methods.
Table 1: Representative AI methods for microbiome research.
Name Key AI architecture Input data Output data Ref
Taxonomic profiling
Metagenome assembly
DeepMAsEd CNN Contigs Contig misassembly probability [113]
ResMiCo ResNet (CNN) Contigs Contig misassembly probability [114]
CheckM2 Hybrid ML MAG genome bins MAG quality [115]
DeepCheck CNN MAG genome bins MAG quality [116]
Metagenome binning
VAMB VAE k-mer composition and abundance Metagenomic bins [117]
CLMB Contrastive learning + VAE Contigs Metagenomic bins [118]
SemiBin Siamese NN Contigs Metagenomic bins [119]
GraphMB VAE + GNN Contigs Metagenomic bins [120]
COMEBin Contrastive multiview NN Contigs Metagenomic bins [121]
Taxonomic classification
DeepMicrobes BiLSTM DNA sequence Taxonomic classification [122]
BERTax DNA BERT DNA sequence Taxonomic classification [123]
Nanopore sequencing basecalling
Chiron CNN + RNN + CTC Raw signal DNA sequence [124]
SACall CNN/ResNet + CTC Raw signal DNA sequence [125]
Mincall CNN/ResNet + CTC Raw signal DNA sequence [126]
Halcyon Inception-CNN + LSTM + attention Raw signal DNA sequence [127]
URnano U-Net (+RNN) Raw signal DNA sequence [128]
Functional annotation & prediction
Gene prediction
Meta-MFDL Deep stacked MLP Metagenomic fragments Metagenomic genes [129]
CNN-MGP CNN ensemble Metagenomic fragments Metagenomic genes [130]
Antibiotic resistance genes identification
DeepARG MLP Metagenomic sequences ARG categories [131]
HMD-ARG CNN Metagenomic sequences ARG categories [132]
HyperVR Ensemble DL Metagenomic sequences ARG categories [133]
ARGNet Autoencoder + CNN (hybrid) Metagenomic sequences ARG categories [134]
ARG-BERT Protein BERT Metagenomic sequences ARG categories [135]
Identification of mobile genetic elements
PlasFlow MLP (plasmid ID) Contigs Plasmid sequences [136]
Deeplasmid LSTM + FC DNA sequences Plasmid sequences [137]
plASgraph2 GNN Assembly graph Plasmid contigs [138]
PPR-Meta Bi-path CNN Metagenomic assemblies Plasmid fragments [139]
Biosynthetic gene clusters prediction
DeepBGC Pfam2vec + BiLSTM Raw sequences BGC classes [140]
e-DeepBGC BiLSTM + enriched embeddings Raw sequences BGC classes [141]
BiGCARP CNN masked LM Raw sequences BGC classes [142]
BGC-Prophet Transformer LM Raw sequences BGC classes [143]
16S rRNA copy number prediction
ANNA16 Stacked MLP + SVM 16S gene sequences 16S rRNA copy number [144]
Mutation/evolution prediction
EVEscape VAE + biophysics Historical sequences Mutation probability [145]
Microbe–X interactions
Microbe-microbe interactions
MicrobialCommunities Random forest Microbial interaction networks Predicted interaction edges [146]
Microbe-host interactions
PIPR Recurrent CNN Protein sequences PPI [147]
DeepPPISP TextCNN Protein sequences PPI [148]
DeepInterface 3D CNN Protein sequences PPI [149]
MaSIF CNN Protein sequences PPI [150]
PECAN Graph CNN Protein sequences PPI [151]
Microbe-disease associations
NinimHMDA GCN MDA database MDA [152]
LGRSH Node2vec MDA database MDA [153]
BPNNHMDA MLP MDA database MDA [154]
MICAH Deep matrix factorization MDA database MDA [155]
Microbe-drug associations
GARFMDA Graph attention network Microbe-drug association Microbe-drug association [156]
GCNATMDA GCN + Graph attention network Microbe-drug association Microbe-drug association [157]
LCASPMDA Graph attention network Microbe-drug association Microbe-drug association [158]
MCHAN GCN with attention Microbe-drug association Microbe-drug association [159]
MDSVDNV Node2vec Microbe-drug association Microbe-drug association [160]
NMGMDA Graph attention network Microbe-drug association Microbe-drug association [161]
OGNNMDA Graph attention network Microbe-drug association Microbe-drug association [162]
STNMDA Structure-Aware Transformer Microbe-drug association Microbe-drug association [163]
Microbial ecology
Microbial composition prediction
cNODE Neural ODE Species collection Microbial composition [164]
MicrobeGNN GNN Bacterial genome Microbial composition [165]
Keystone species identification
DKI Neural ODE Microbial composition Keystoneness [166]
Microbial dynamics prediction
MiRNN Tailored RNN Dynamic community data Metabolite concentrations [167]
MTSF-DG GNN Historical timeseries Future timeseries [168]
Control of microbial communities
ROCC RL Current microbial population levels Optimal feed actions [169]
Microbiome data simulation, augmentation, and imputation
MB-GAN GAN Microbial abundances Microbial abundances [170]
mbSparse Autoencoder + CVAE Raw taxonomic count matrix Imputed taxonomic count matrix [171]
Microbial source tracking
SourceTracker Gibbs sampling Taxonomic profile Source contributions [172]
FEAST Fast expectation-maximization Taxonomic profile Source contributions [173]
STENSL Expectation-maximization Taxonomic profile Source contributions [174]
ONN4MST Ontology-aware Neural Network (ONN) Taxonomic profile Source contributions [175]
Metabolic Modeling
Gap filling: inferring missing reactions
AGORA2 GEM Microbial genome sequences Metabolic capabilities [176]
CHESHIRE GCN Metabolic networks Candidate reactions [177]
Retrosynthesis: breaking down a target molecule
RetroPath RL RL Target compound, reaction-rule set Predicted biosynthetic pathways [178]
Model microbial metabolism
SPAM-DFBA RL Current community statet Flux-regulation actions [179]
Precision nutrition
Nutrition profile correction
METRIC MLP Gut microbiome, noised dietary profiles Corrected dietary profiles [180]
Metabolomic profile prediction
MiMeNet MLP Microbiome composition Metabolome composition [181]
MelonnPan linear regression Microbiome composition Metabolome composition [182]
NED MLP Microbiome composition Metabolome composition [183]
mNODE neural ODE Microbiome composition Metabolome composition [184]
Personalized diet recommendation
McMLP MLP Baseline microbial compositions, dietary intervention strategy, baseline metabolomic profiles Metabolomic profiles [185]
Clinical Microbiology
Microorganism detection, identification and quantification
BIOINTEL CNN Microscopy imaging Bacteria type prediction [186]
Antimicrobial susceptibility evaluation
ANN MLP Mutation patterns Resistance to the protease inhibitor lopinavir [187]
GoPhAST-R RF NanoString data Probability of resistance to antibiotics [188]
CART Classification and Regression Trees Pyrazinamide MIC MIC threshold [189]
Disease diagnosis, classification, and clinical outcome prediction
Ph-CNN Phylogeny-based CNN Taxonomic profile Sample categories  [190]
PopPhy-CNN Phylogeny-based + abundance + CNN Taxonomic profile Sample categories [191]
taxoNN Phylogeny cluster-based CNN Taxonomic profile Sample categories [192]
MDeep Phylogeny-based CNN Taxonomic profile Sample categories [193]
PM-CNN Phylogeny-based CNN Taxonomic profile Sample categories  [194]
DeepMicro Autoencoders + MLP Taxonomic profile Sample categories [122]
GDmicro GCN Taxonomic profile Sample categories [195]
MLM Transformer Taxonomic profile Sample categories [196]
MSFT-transformer Multistage fusion tabular Transformer Taxonomic and functional profiles Sample categories [197]
Integration of various feature types
MDL4Microbiome MLP Taxonomic and functional profiles Sample categories [198]
Prevention and Therapeutics
Vaccine design
Vaxi-ML XGBoost Protein features Protective-antigen prediction [199]
Vaxi-DL MLP Protein sequences Protective-antigen prediction [200]
Probiotic mining
iProbiotics SVM Whole-genome sequences Probiotic property prediction per genome [201]
metaProbiotics Random forest Metagenomic bins Per-bin probiotic score [202]
Antibiotic discovery
MPNN GCN Molecular graphs Predicted molecular properties [203]
Chemprop DL GCN + MLP Molecular graphs Predicted molecular properties [204]
Antimicrobial peptides identification & design
MLBP CNN Peptide sequence Peptide categories [205]
AMPGANv2 cGAN Peptide sequence Potent AMPs [206]
EvoGradient CNN, LSTM, Transformer Peptide sequence Potent AMPs [207]
deepAMP CVAE Peptide sequence Potent AMPs [208]
Phage therapy
INHERIT DNA BERT DNA contigs Phage vs. non-phage classification [209]
PHACTS RF Proteome profiles of bacteriophages Phage lifestyle classification [210]
DeePhage CNN DNA fragments Phage lifestyle classification [211]
DeePhafier Self-attention Phage protein features Phage lifestyle classification [212]
PhaTYP Transfomer Phage genome Phage lifestyle classification [213]
PHIAF GAN Phage and host genome Predicted phage-host interaction [214]
PHPGAT GAT Integrated phage-phage, host-host and phage-host links Predicted phage-host interaction [215]
SpikeHunter ESM-2 Metagenomic contigs Phage categories [216]
PhANNs MLP Phage protein sequences Protein structural class [217]
VirionFinder CNN Viral protein sequences Virion vs. non-virion classification [218]
DeePVP CNN Phage protein sequences Virion vs. non-virion classification [219]
PhaVIP Transformer Phage protein sequences Virion vs. non-virion classification [220]
ESM-PVP ESM-2 Phage protein sequences Virion vs. non-virion classification [221]
DeepLysin Ensemble model Phage protein sequences Antibacterial activity score per lysin [222]
DeepMineLys CNN Lysins protein sequences Predicted phage lysins [223]
@@ 3.1 Taxonomic Profiling
Figure 1: Application scenarios of AI in microbiology and microbiome research. The application scenarios of AI in microbiology and microbiome research are diverse, spanning (1) taxonomic profiling (e.g., metagenome assembly, binning, taxonomic classification, and nanopore sequencing basecalling), (2) functional annotation and prediction (including gene prediction, antibiotic resistance gene and plasmid identification, biosynthetic gene clusters prediction, 16S rRNA copy number prediction and mutation/evolution prediction), (3) microbe-host/disease/drug interactions, (4) microbial ecology (such as microbial interaction networks, composition and colonization outcome predictions and keystone species identification), (5) metabolic modeling, (6) precision nutrition (e.g., nutrition profile correction, metabolomic profile prediction and personalized diet recommendation), (7) clinical microbiology (such as microorganism detection, identification and quantification, antimicrobial susceptibility evaluation and disease diagnosis, classification, and clinical outcome prediction), and (8) applications in prevention and therapeutics (including peptides identification and generation, probiotic mining, antibiotic discovery, phage therapy and vaccine design).
A fundamental goal of microbiology and microbiome research is determining the compositions of microbial communities, i.e., identifying and quantifying different types of microorganisms (such as bacteria, fungi, viruses, and archaea) present in a given sample. This involves analyzing their relative abundances and diversity, often using DNA sequencing techniques. Currently, three generations of DNA sequencing techniques are available for microbiome research. The first-generation sequencing utilizes the chain termination method, offering read lengths of 500-1000 base pairs [224]. Second-generation sequencing, also known as next-generation sequencing (NGS), includes methods such as pyrosequencing, sequencing by synthesis, and sequencing by ligation, with read lengths ranging between 50 and 500 bp [225, 226, 227]. Two key NGS applications in microbiome research are (1) amplicon sequencing, which targets small fragments of one or two hypervariable regions of the 16S rRNA gene (for archaea and bacteria) or 18S rRNA gene (for fungi); and (2) metagenomic shotgun sequencing, which comprehensively samples all genes in all organisms present in a given community. The third-generation sequencing performs single-molecule sequencing, offering long reads with lengths reaching tens of kilobases on average [228]. In the following, we discuss applications of AI techniques in various aspects of taxonomic profiling.
@@ 3.1.1 Metagenome assembly
Metagenomics refers to the direct study of the entire genomic information contained in a microbial community. Metagenomics avoids isolating and culturing individual microorganisms in a community and provides a way to study microorganisms that cannot be isolated and cultured. There are two main approaches for processing metagenomic sequencing data: (1) assembly-based and (2) reference database-based. The goal of the assembly-based approach is to construct and annotate the so-called metagenome-assembled genomes (MAGs) [229]. The construction of MAGs includes two main steps: assembly and binning. Assembly refers to the process of reconstructing longer sequences (contigs) from DNA sequencing reads, which can vary in length and type depending on the technology used. This involves piecing together overlapping reads—using short-read, linked-read, long-read, or hybrid assemblers—to form contiguous sequences that represent parts of the genomes present in the microbial community [230, 231].
@@
Machine learning has been widely used in the quality control of metagenomic assembly. Many factors (e.g., sequencing errors, variable coverage, repetitive genomic regions, etc.) can produce misassemblies. For taxonomically novel genomic data, detecting misassemblies is very challenging due to the lack of closely related reference genomes. Machine learning methods can identify misassembled contigs in a reference-free manner. Representative methods include DeepMAsEd [113], ResMiCo [114] CheckM2 [115], and DeepCheck [116]. DeepMAsEd is based on CNN. Denote a contig as a sequence of nucleotides. At each position in the sequence, the concatenation of two types of information (raw sequence and read-count features) yields the input vector. To train and evaluate DeepMAsEd, one can generate a synthetic dataset of contigs, read counts, and binary assembly quality labels. As an extension of DeepMAsEd, ResMiCo is based on ResNet, a variant of CNN. The key feature of ResNet is the introduction of skip connections, which effectively solves the degradation problem of deep neural networks [232]. Compared to DeepMAsEd, ResMiCo leveraged a much more informative input vector computed from raw reads and contigs. CheckM2 is a machine learning framework to predict genome quality of MAGs [115]. CheckM2 uses both CNN and the gradient boosted (GB) decision tree for completeness prediction, and uses the GB model to predict contamination. Since CheckM2 employs two separate models to predict the completeness and contamination of MAGs, it may overlook the inherent correlation between these two tasks. DeepCheck is a multitasking deep learning framework based on ResNet and the attention mechanism to simultaneously predict MAG completeness and contamination [116]. We notice that those methods typically use a carefully designed input vector. For example, ResMiCO transforms raw reads and contigs into carefully selected positional features, such as coverage (number of reads aligned to position), number of single-nucleotide variants (SNVs), mean alignment score, etc. CheckM2 and DeepCheck calculate feature vectors from the genome properties of synthetic MAGs, e.g., the genome length, number of coding sequences and individual amino acid counts, as well as annotation of predicted proteins using KEGG (the Kyoto Encyclopedia of Genes and Genomes). It would be interesting to explore if we can use a more advanced deep learning architecture (e.g., the Transformer) or a hybrid learning approach (e.g., CNN + RNN) to directly deal with the raw sequence data, avoiding the manual design of the input vector.
@@ 3.1.2 Metagenome binning
Metagenomic binning involves grouping those assembled sequences into clusters (bins or MAGs) that correspond to different species or genomes [233]. Metagenomic binning helps us identify and categorize different microorganisms present in a metagenomic sample, even if they are not fully assembled into complete genomes. There are many methods for metagenomic binning [234, 235, 236, 237]. Several binning methods are based on deep learning, e.g., VAMB [117], CLMB [118], SemiBin [119], GraphMB [120], and COMEBin [121]. VAMB (Variational Autoencoders for Metagenomic Binning) uses VAE to encode sequence coabundance and $k$-mer distribution information, and clusters the resulting latent representation into genome clusters and sample-specific bins [117]. As an extension of VAMB, CLMB (Contrastive Learning framework for Metagenome Binning) can efficiently eliminate the disturbance of noise and produce more stable and robust results [118]. CLMB is based on contrastive learning, an machine learning approach that focuses on extracting meaningful representations by contrasting positive and negative instances [118]. SemiBin employs deep siamese neural networks to exploit the information in reference genomes, while retaining the capability of reconstructing high-quality bins that are outside the reference dataset [119]. Here, a siamese neural network (a.k.a. twin neural network) is a neural network that uses the same weights while working in tandem on two different input vectors to compute comparable output vectors [238]. GraphMB integrates GCN with assembly graphs to improve binning accuracy [120]. It models each contig using VAE for feature generation and aggregates these features using a GCN. This method accounts for read coverage in its loss function and uses iterative medoid clustering to finalize the binning. COMEBin is the latest metagenomic binning method [121]. This method is based on contrastive multiview representation learning. It introduces a data augmentation approach that generates multiple views for each contig, enabling contrastive learning and yielding high-quality representations of the heterogeneous features. Moreover, it incorporates a “Coverage module” to obtain fixed-dimensional coverage embeddings, which enhances its performance across datasets with varying numbers of sequencing samples. It also adapts an advanced community detection algorithm, Leiden, specifically for the binning task, considering single-copy gene information and contig length. COMEBin outperformed VAME and SemiBin on various simulated and real datasets, especially in recovering near-complete genomes from real environmental samples.
@@ 3.1.4 Nanopore sequencing basecalling
Since its introduction, nanopore sequencing has revolutionized microbial community studies by enabling portable, real-time long-read sequencing (with reads longer than a few thousand bases) at low cost [257, 258]. Its long reads overcome key challenges in reconstructing complex genomic structures and assembling high-quality MAGs. Coupled with rapid sample-to-answer workflows and streamlined analysis pipelines, nanopore technology now offers an accessible and powerful platform for real-time, in-field monitoring of environmental microbiomes. The basic principle of nanopore sequencing is to pass an ionic current through a nanopore and measure the change in current when a biomolecule passes through or approaches the nanopore. Information about the change in current can be used to identify the molecule, a process often referred to as basecalling. There are two challenges in basecalling. First, the current signal level is most dominantly influenced by the several nucleotides that reside inside the pore at any given time, rather than a single base. Second, DNA molecules do not translocate at a constant speed. Basecalling is conceptually similar to speech recognition. Both processes involve interpreting complex signals to extract meaningful sequences-DNA bases in the case of basecalling, and spoken words in the case of speech recognition. Much like the evolution of speech recognition methods, computational methods for basecalling have evolved from statistical tests to hidden Markov models and finally deep learning models. Those methods are often referred to as basecallers.
Various deep learning models have been developed for basecalling. Chiron is the first deep learning model that can translate raw electrical signal directly to nucleotide sequence [124]. It applied a CNN to extract features from the raw signal, an RNN to relate such features in a temporal manner, and a connectionist temporal classification (CTC) decoder to create the nucleotide sequence. Here, CTC enabled us to generate a variant length base sequence for a fixed-length signal window through output-space searching, avoiding explicit segmentation for basecalling from raw signals. Similar to the Chiron architecture, SACall [125] integrated CNN with Transformer (Lite Transformer or LSTM) and CTC. Similar approaches are also implemented in CATCaller [259] and Bonito [260]. Mincall [126] (or Causalcall [261]) directly integrated ResNet (or causal dilated CNN) with CTC. Halcyon used a different architecture  [127]. It combines a novel inception-block-based CNN module, an LSTM-based encoder, and an LSTM-based decoder using an attention mechanism. The inception-block-based CNN module aims to extract local features of input raw signal and reduce the dimension of the input timestep axis. The LSTM-based encoder captures long-time dependencies in the timestep dimension and deals with the variable lengths of inputs. The attention mechanism allows the decoder to focus on specific parts of the input sequence when generating each element of the output sequence.